wHAT ARE THE STEPS TO DRAWING A PIE CHART.....HOW DO YOU DRAW A PIE CHART OR CONSTRUCT A PIE CHART CAN YOU EXPLAIN THE STEPS
Found 2 solutions by stanbon, venugopalramana:
Answer by stanbon(57984)

Let's assume your data is 20,30,40
Convert each value to a percentage:
20/90= 22%
30/90= 33%
40/90= 44%
Draw a circle.
Mark a pie-shaped sector = 44%
Mark a pie-shaped sector = 33%
Mark a pie-shaped sector = 22%
That is your pie chart.
Cheers.
Stan H.

Answer by venugopalramana(3286)

wHAT ARE THE STEPS TO DRAWING A PIE CHART.....HOW DO YOU DRAW A PIE CHART OR CONSTRUCT A PIE CHART CAN YOU EXPLAIN THE STEPS
LET US TAKE THE FOLLOWING DATA ON GRADES OBTAINED BY STUDENTS IN A CLASS
A = 50
B= 100
C=50
TOTAL=200
MAKE A TABLE OF % .
A = 50*100/200=25
B=100*100/200=50
C=50*100/200=25
TOTAL = 100
NOW TAKING TOTAL AS 360 DEGREES CONVERT THE ABOVE PERCENTAGES IN TO DEGREES
A=360*25/100=90
B=360*50/100=180
C=360*25/100=90
TOTAL = 360 DEGREES.
NOW DRAW ANY CIRCLE TAKING A CONVENIENT RADIUS WITH O AS CENTER .
DRAW ANY LINE OP SAY WITH P ON THE CIRCUMFERENCE OF THE CIRCLE.SUPPOSE WE DRAW
IT HORIZONTALLY.
NOW TO SHOW A GRADE FROM OP MEASURE AT 'O' 90 DEGREES CORRESPONDING TO 90 DEGREES WE GOT FOR 'A' GRADE STUDENTS.. TO GET LINE OA WITH A ON CIRCUMFERENCE,SUCH THAT ANGLE AOP=90 DEG.
NOW TO SHOW B GRADE FROM OA MEASURE AT 'O' 180 DEGREES CORRESPONDING TO 180 DEGREES WE GOT FOR 'B' GRADE STUDENTS.. TO GET LINE OB WITH B ON CIRCUMFERENCE,SUCH THAT ANGLE AOB=180 DEG.
NOW TO SHOW C GRADE FROM OB MEASURE AT 'O' 90 DEGREES CORRESPONDING TO 90 DEGREES WE GOT FOR 'C' GRADE STUDENTS.. TO GET LINE OC WITH C ON CIRCUMFERENCE,SUCH THAT ANGLE BOC=90 DEG.
WHEN IT IS THE END OF PLOT , YOU WILL NOTICE THAT C COINCIDES WITH 'P'.
IF NOT ,EITHER PLOTTING IS NOT COMPLETE OR YOUR CALCULATION IS WRONG.
see the drawing below.
now shade different zones IN DIFFERENT COLORS OR HATCHING
POA , LABELING IT AS A GRADE
AOB,LABELING IT AS B GRADE
AND FINALLY
BOP(BOC...C COINCIDES WITH P),LABELING IT AS C GRADE